# Brain functions emerge as thermal equilibrium states of the connectome

### v2

@@ Results
6. Pure functional connectomes from mixed states. Now notice that mixed functional states given by Eq. (4) can also be expressed as a matrix product via
@@
where $[\operatorname{{\bf x}}^{\bullet|\beta}]$ is the $N\times N$-matrix whose columns are the $\operatorname{NEP_{\beta}}$ vectors $\operatorname{{\bf x}}^{v|\beta}$ and $\operatorname{{\bf p}}=(\operatorname{{\bf p}}_{v})_{v}$ is a distribution over neurons. Since each column of $[\operatorname{{\bf x}}^{\bullet|\beta}]$ measures potential functional connections from a neuron to the other neurons, the whole matrix measures potential functional connections between each pair of neurons in the connectome. Thus, a mixed $\beta$–KMS state $\operatorname{{\bf x}}$ encodes 1) the potential functional connections between each neuron pair at inverse temperature $\beta$, and 2) the expected value of the upstream functional connectivity onto each neuron given the distribution $\operatorname{{\bf p}}$ on the neurons. It follows that for each $\beta$ value, $[\operatorname{{\bf x}}^{\bullet|\beta}]$ is a connectivity matrix encoding the substrates underlying all functional states of the connectome.
@@
Moreover, using the notion of PFCs introduced in the previous paragraph, we get a pure functional connectivity matrix at inverse temperature $\beta$ by restricting $[\operatorname{{\bf x}}^{\bullet|\beta}]$ to its statistically significant values. The resulting matrix is the adjacency matrix of a weighted directed network $\mathcal{F}^{\beta}$ whose edges the PFCs at the given $\beta$ value, and for this reason we call it the ’pure functional connectome’ at inverse temperature $\beta$. Specifically, $\mathcal{F}^{\beta}$ represents potential functional connections predicted by the topology of the anatomical network.
We investigated $\mathcal{F}^{\beta}$ at the functional $\beta_{f}=4.51$ (see Supplementary Data 3 for the edgelist) and found that it reveals connectivity patterns that coincide with circuitries that have been reported by researchers to be the functional substrates of well studied C. elegans complex behaviors such as locomotion and touch-induced movement [63, 12], indicating that our network model provides a theoretical basis for understanding the functional organization of a nervous system.
Figure 4: Pure functional connectivity matrix. Representation of part of the adjacency matrix of the pure functional connectome $\mathcal{F}^{\beta}$ at $\beta=4.51$. Neurons on the x-axis are the ones with the largest in-degrees in $\mathcal{F}^{\beta}$, and on the y-axis are those with the highest out-degrees.
7. Centrality of the locomotory circuit. Indeed, $\mathcal{F}^{\beta}$ has 8932 weighted directed edges, compared to the network $\mathcal{K}$ formed by the structural connectivity of all neurons which has 4927 weighted directed edges (see Supplementary Data 2). It represents $11.68\%$ of all the potential functional connections given by the adjacency matrix $[\operatorname{{\bf x}}^{\bullet|\beta}]$, implying that only about $12\%$ of all neural $\beta$–emittance connections are PFCs. This network is schematically represented in Figure 3.
In-degree ranking
$\mathcal{F}^{\beta}$ $\mathcal{K}$
Neuron $k^{in}$ $w^{in}$ $k^{in}$ $w^{in}$
AS08 1 10 278 278
DA06 2 2 248 248
AS07 3 6 269 270
DA07 4 44 256 275
VA08 5 1 193 115
DB05 6 72 236 274
VA10 7 30 257 259
DB06 8 32 249 273
AS09 9 15 270 256
VA11 10 8 183 179
DA04 11 11 142 202
VA06 12 21 181 217
AS06 13 94 223 255
AS05 14 87 208 257
DA05 15 16 92 134
VA07 16 85 222 216
AS04 17 178 235 272
AS10 18 105 179 234
VA05 19 35 154 162
VA09 20 24 182 164
DA08 21 9 82 85
AS02 22 147 207 254
DA03 23 23 98 143
VB06 24 173 226 263
VA04 25 79 156 168
VB09 26 38 211 181
AS03 27 130 192 253
AS01 28 101 178 243
LUAR 29 84 170 204
VA03 30 62 153 136
SABVR 31 144 210 242
PVCL 32 5 8 8
VD08 33 109 273 123
DA02 34 76 66 118
VB11 35 127 194 245
VB07 36 194 227 262
SABVL 37 97 184 214
DA01 38 57 155 62
VB10 39 177 239 209
PHCR 40 185 212 231
AS11 41 74 74 147
PVCR 42 7 9 6
AVAR 43 4 2 2
AVAL 44 3 1 1
Out-degree ranking
$\mathcal{F}^{\beta}$ $\mathcal{K}$
Neuron $k^{out}$ $w^{out}$ $k^{out}$ $w^{out}$
PLML 1 259 272 158
DA07 2 273 274 146
PVCR 3 113 5 163
SIBDR 4 52 248 74
PVCL 5 185 4 115
VB07 6 7 261 200
SABVL 7 197 269 198
SIBDL 8 266 224 209
DB05 9 1 276 150
SIBVL 10 137 176 208
AS08 11 183 279 143
DB06 12 228 275 147
AS10 13 196 253 83
SIAVR 14 216 250 187
VA10 15 231 255 196
SIADR 16 136 178 31
SIBVR 17 138 181 57
VD11 18 257 244 109
VD05 19 84 217 216
AVAR 20 18 1 268
AVAL 21 2 2 114
SABVR 22 34 268 108
ALA 23 17 179 55
AS09 24 88 273 144
DB07 25 112 201 191
PVR 26 246 8 117
ALMR 27 247 183 121
SABD 28 33 259 195
AVBR 29 4 6 26
DVA 30 129 3 278
VC04 31 169 245 203
VD08 32 93 271 156
VD06 33 175 238 82
AS07 34 3 267 190
DA06 35 20 254 111
AVBL 36 189 7 13
SIADL 37 51 235 212
PVWL 38 254 222 86
AVDR 39 5 9 7
PVNR 40 269 11 101
VD10 41 256 262 153
AVDL 42 19 18 164
DVC 43 58 29 51
BDUL 44 16 173 29
We observed a substantial topological discordance between $\mathcal{F}^{\beta}$ and $\mathcal{K}$. For instance, the motor neuron classes AS, VA, DA, VB, and DB, have the highest in-degree centrality $\mathcal{F}$, with AS08, DA06, AS07, DA07, VA08, DB05, VA10, DB06, AS09, VA11 being the 10 most central neurons in terms of incoming connections, and specifically AS08, which is ranked 278 in the anatomical network (see Table 2), receiving 219 PFCs. Notably, none of these neurons is among the most in-connected in the structural connectome ($k^{in}=2$ for AS08 and 7 for DA06) which has the command interneuron classes AVA, AVB, AVE, AVD, and PVC receiving the highest numbers of incoming connections. Moreover, VA08 has the highest weighted in-degree $w^{in}$ followed by DA06, and DB05 has the largest weighted out-degree $w^{out}$, while they are poorly connected anatomically. Additionally, some of these neurons are highly ranked in terms of out-degree and weighted out-degree (see Table 2). Specifically, DA07 has the second largest number of outgoing connections ($k^{ou}=64$), compared to its $3$ anatomical weighted connections, VB07 has out-degree rank 6 and weighted out-degree rank 7, and DB05 has out-degree rank 9 and weighted out-degree rank 1. Interestingly, PVCL and PVCR are the most functionally in- and out-connected among the command interneurons most of whom have their anatomical status downgraded in the pure functional connectome. Indeed, on the one hand, AVAL, AVAR, PVCL, and PVCR have their respective structural weighted in-degree ranks almost maintained in $\mathcal{F}^{\beta}$, while their weighted out-degree ranks are not, and on the other, PVCL and PVCR have their high structural out-degree ranks almost maintained in $\mathcal{F}^{\beta}$ while their weighted out-degree ranks are not.
8. Mechanoreceptor neurons have higher functional out-degree. Additionally, we found that the posterior mechanoreceptor neuron [23] PLML is the most central in number of outgoing connections in $\mathcal{F}^{\beta}$ ($k^{out}=66$), and the anterior mechanoreceptor ALMR is among the 30 most out-connected with $k^{out}=54$ in $\mathcal{F}^{\beta}$ (ranked 27 in Table 2). By comparison, both neurons have very low out-degrees of respectively 4 and 12 and low ranks of 272 and 183 in the structural network $\mathcal{K}$, and both neurons have low ranks in terms of weighted out-degree (259 and 247, respectively). More generally, many mechanoreceptor neurons [31, 23, 64], including ALA, ASHR, AVM, DVA, FLPL/R, OLQVL, etc., are among the most out-connected in $\mathcal{F}^{\beta}$.
9. Integration capacity of neurons. Let us now consider a particular mixed KMS state defined as follows. Letting $\operatorname{{\bf p}}$ in (7) be the uniform distribution $\operatorname{{\bf p}}_{v}=1/N$ on $V$, the resulting mixed state $\langle\operatorname{{\bf x}}^{v|\beta}\rangle=\frac{1}{N}\sum\operatorname{{\bf x}}^{v|\beta}$ is the mean $\operatorname{NEP_{\beta}}$, and the probability of a neuron $u$ to receive signals in this state is the average neural $\beta$–emittance to $u$ at inverse.
@@
Specifically, it is the expected relative weight of all possible PFCs onto the neuron, excluding its neural self-emittance at inverse temperature $\beta$. A high $\operatorname{\bf IC}$ value at a given $\beta$ means a high likelihood of receiving any information flowing through the network, reflecting two topological features of the neuron: 1) it is the target of a relatively large number of independent flow pathways of certain lengths; in other words, a high capacity for the neuron to integrate information flow coming from many different neurons that do not belong to the same cycles; in particular, the $\operatorname{\bf IC}$ of a highly connected neuron with nearly no closed paths will be close to 1, whereas an equally connected neuron whose almost all incoming paths are cycles will have an $\operatorname{\bf IC}$ close to 00; and 2) it has substantial redundancies among its incoming pathways; thus, the neuron is well-situated in the connectome to receive inputs from multiple neurons via a number of parallel pathways of certain range.
Table 3: Classification of neurons by IC
Group 1 ADER, AS01, AS02, AQR, ALNR, ALMR, AS03, AS04, AS05, AS06, AS07, AS08, AS09, AS10, AS11, ASHR, AVAL, AVAR, AVBR, AVBL, AVDL, AVDR, AVEL, AVER, AVG, AVHL, AVHR, AVJL, AVJR, AVKL, AVL, AVM, DA01, DA02, DA03, DA04, DA05, DA06, DA07, DA08, DA09, DB01, DB02, DB03, DB04, DB05, DB06, DB07, DD01, DD02, DD03, DD04, DD05, DD06, DVA, DVB, DVC, FLPL, FLPR, LUAL, LUAR, PDA, PDB, PDEL, PDER, PHBL, PHCL, PHCR, PLML, PLMR, PQR, PVCL, PVCR, PVDL, PVDR, PVM, PVNL, PVNR, PVPL, PVPR, PVR, PVT, PVWL, PVWR, RID, RIMR, SAAVR, SABD, SABVL, SABVR, SDQL, SDQR, SIBDL, SIBVL, URYDL, URYDR, URYVL, URYVR, VA01, VA02, VA03, VA04, VA05, VA06, VA07, VA08, VA09, VA10, VA11, VA12, VB01, VB02, VB03, VB04, VB05, VB06, VB07, VB08, VB09, VB10, VB11, VC01, VC06, VD01, VD02, VD03, VD04, VD05, VD06, VD07, VD08, VD09, VD10, VD11, VD12, VD13
Group 2 ADLR, ADFR, ADFL, AFDL, AFDR, ADEL, ADLL, AIAL, AIAR, AIBL, AIBR, AIML, AIMR, AINL, AINR, AIYL, AIYR, AIZL, AIZR, ALML, ALNL, ASEL, ASER, ASGL, ASGR, ASHL, ASIL, ASIR, ASJL, ASJR, ASKL, ASKR, AUAL, AUAR, AWAL, AWAR, AWBL, AWBR, AWCL, AWCR, BAGL, BAGR, CEPDL, CEPDR, CEPVL, CEPVR, HSNL, IL1DL, IL1DR, IL1L, IL1R, IL1VL, IL1VR, IL2DL, IL2DR, IL2L, IL2R, IL2VL, IL2VR, OLLL, OLLR, OLQDL, OLQDR, OLQVL, OLQVR, RIAL, RIAR, RIBL, RIBR, RICL, RICR, RIH, RIPL, RIPR, RIVL, RIVR, RMDDL, RMDDR, RMDL, RMDR, RMDVL, RMDVR, RMED, RMEL, RMER, RMEV, RMGL, RMGR, RMHL, RMHR, SAADL, SAADR, SAAVL, SIADL, SIADR, SIAVL, SIAVR, SIBDR, SMBDL, SMBDR, SMBVL, SMDDL, SMDDR, SMDVL, SMDVR, URADL, URADR, URAVL, URAVR, URBL, URBR, URXL, URXR, VC02, VC03, VC04, VC05
Group 3 ADAR, ALA, AVFL, AVFR, AVKR, BDUL, BDUR, PHAL, PHAR, PHBR, RIGL, RIS, RMFL, RMFR, SIBVR, SMBVR
Group 4 ADAL, HSNR, PLNL, PLNR, PVQL, PVQR, RIFL, RIFR, RIGR, RIML, RIR
@@
In one sense, neurons whose $\operatorname{\bf IC}$ is high at small temperature values but low at larger ones are structurally positioned to mediate relatively short-range communications, making them ’local integrators’ [43], whereas those with high $\operatorname{\bf IC}$ at higher temperatures are ’global integrators’ capable of mediating multiple long-range incoming communications.To better understand the functional properties of the $\operatorname{\bf IC}$, we analyzed its variation with respect to temperature. We observed four main neuron groups showing different integration behaviors (see Table 3 for the complete classification) schematically illustrated by Fig. 5 and characterized as follows.

### v3

@@ II.3 Dependence on network topology
Now notice that mixed functional states given by Eq. (4) can also be expressed as a matrix product via
@@
where $\operatorname{{\bf x}}^{\bullet|\beta}]$ is the $N\times N$-matrix whose columns are the $\operatorname{NEP_{\beta}}$ vectors $\operatorname{{\bf x}}^{v|\beta}$ and $\operatorname{{\bf p}}=(\operatorname{{\bf p}}_{v})_{v}$ is a distribution over neurons. Since each column of $\operatorname{{\bf x}}^{\bullet|\beta}]$ measures potential functional connections from a neuron to the other neurons, the whole matrix measures potential functional connections between each pair of neurons in the connectome. Thus, a mixed $\beta$–KMS state $\operatorname{{\bf x}}$ encodes 1) the potential functional connections between each neuron pair at inverse temperature $\beta$, and 2) the expected value of the upstream functional connectivity onto each neuron given the distribution $\operatorname{{\bf p}}$ on the neurons. It follows that for each $\beta$ value, $\operatorname{{\bf x}}^{\bullet|\beta}]$ is a connectivity matrix encoding the substrates underlying all functional states of the connectome.
@@
Moreover, using the notion of PTFCs introduced in the previous paragraph, we get a purely topological functional connectivity matrix at inverse temperature $\beta$ by restricting $\operatorname{{\bf x}}^{\bullet|\beta}]$ to its statistically significant values. The resulting matrix is the adjacency matrix of a weighted directed network $\mathcal{F}^{\beta}$ whose edges the PTFCs at the given $\beta$ value, and for this reason we call it the ’purely topological functional connectome’ at inverse temperature $\beta$. Specifically, $\mathcal{F}^{\beta}$ represents potential functional connections predicted by the topology of the anatomical network.
II.5 Integration capacity
Let us now consider a particular mixed KMS state defined as follows. Letting $\operatorname{{\bf p}}$ in (7) be the uniform distribution $\operatorname{{\bf p}}_{v}=1/N$ on $V$, the resulting mixed state $\langle\operatorname{{\bf x}}^{v|\beta}\rangle=\frac{1}{N}\sum\operatorname{{\bf x}}^{v|\beta}$ is the mean $\operatorname{NEP_{\beta}}$, and the probability of a neuron $u$ to receive signals in this state is the average neural $\beta$–emittance to $u$ at inverse.
@@ II.5 Integration capacity
Specifically, it is the expected relative weight of all possible PTFCs onto the neuron, excluding its neural self-emittance at inverse temperature $\beta$. A high $\operatorname{\bf IC}$ value at a given $\beta$ means a high likelihood of receiving any information flowing through the network, reflecting two topological features of the neuron: 1) it is the target of a relatively large number of independent flow pathways of certain lengths; in other words, a high capacity for the neuron to integrate information flow coming from many different neurons that do not belong to the same cycles; in particular, the $\operatorname{\bf IC}$ of a highly connected neuron with nearly no closed paths will be close to 1, whereas an equally connected neuron whose almost all incoming paths are cycles will have an $\operatorname{\bf IC}$ close to 0; and 2) it has substantial redundancies among its incoming pathways; thus, the neuron is well-situated in the connectome to receive inputs from multiple neurons via a number of parallel pathways of certain range.
@@
In one sense, neurons whose $\operatorname{\bf IC}$ is high at small temperature values but low at larger ones are structurally positioned to mediate relatively short-range communications, making them ’local integrators’ [21], whereas those with high $\operatorname{\bf IC}$ at higher temperatures are ’global integrators’ capable of mediating multiple long-range incoming communications.
III Results
Figure 1: Connectivity states of AFDR. Schematic visualizations of different types of connectivity states AFDR: structural connectivity (gray arrows), and the emittance networks (green and purple arrows) obtained from neural $\beta$-emittance profiles ($\operatorname{NEP_{\beta}}$) at different values of inverse temperature $\beta>\beta_{c}$, where $\beta_{c}$ is the critical inverse temperature. a: the structural connectivity of AFDR; the weights of the connections are the components of the structural connectivity vector $\operatorname{{\bf k}}^{AFDR}$. b: At $\beta_{s}=2.5\beta_{c}$, the emittance network coincides with the structural connectivity of AFDR. However, the connection to RMDVR (represented by the purple arrow) is not statistically significant ($p\sim 0.2$), and therefore not a purely topological functional connection (PTFC). In other words, when the system is at this $\beta_{s}$–KMS state, AFDR can functionally communicates only with its direct neighbors at the same probability distribution as its structural connectivity state. c: At a relatively higher temperature $1/\beta$, long-range communications emerge (i.e. neurons that are not post-synaptic to AFDR, here positioned at the external concentric circle) while the neuron maintains neural emittances onto its direct neighbors (neurons positioned at the internal circle), albeit with lower intensity. d: At a very high temperature $1/\beta$ with $\beta$ close to the critical value, more longer range connections emerge, while some short range ones are either no longer present or not statistically significant. In both c and d, only the PTFCs ($p<0.05$) are considered. Greens are sensory neurons, yellows are interneurons, and pinks are motor neurons.
III.1 Emittance networks
To illustrate our model, we reprent in Figs.1b–d the emittance networks of the thermosensory neuron $v={\rm AFDR}$ at different values of $\beta$, illustrating how they evolve as the (inverse) temperature varies. We see that increasing the temperature allows the emergence of long-range emittance connections which are facilitated by large amounts of parallel pathways of different lengths. Indeed, it is an immediate consequence of formula (3) that parallel paths and redundancies are critical to the emergence of strong neural emittances, and the lengths of pathways and the degree of redundancies that significantly contribute to their weights depend on temperature. More precisely, lower temperatures (i.e., higher values of $\beta$) favor parallel direct anatomical connections and short paths. Whereas, higher temperatures enhance the contributions of redundant long pathways of different lengths while lessening the importance of direct anatomical connections in the emittance weights. For example, AFDR has 13 synaptic connections and 52 anatomical paths of length 2 onto the interneuron AIYR, contributing greatly to the important weight of its neural emittance onto the latter at low temperature (Fig.1b). And at higher temperature values, while the neural emittance of AFRD $\longrightarrow$ AIYR decreases, other connections are ’established’ with neurons onto which AFDR can only be connected through thousands or millions of parallel anatomical paths of length $\geq 2$. For instance, there are 117 paths of length 2 and about 4 millions of length between 3 and 5 linking AFDR to RIAR, contributing to the neural emittance onto this neuron at $\beta=1.05\beta_{c}$ (Fig.1c), with $\beta_{c}=4.2958$. Similarly, the neural emittance to the ventral cord motor neuron VA11 that appears at $\beta=1.01\beta_{c}$ (Fig.1d) is largely due to the more than its 230 million incoming paths of length $\geq 5$ from AFDR.
III.2 Structure vs function
Computing the $\operatorname{\bf sfd}$.
We found that $\operatorname{\bf sfd}(v,\beta)\sim 0$ at $\beta=2.5\beta_{c}$ for each of the 280 neurons. Therefore, we have $\beta_{s}\approx 10.7394$. In particular, going back to the example of AFDR, its neural $\beta_{s}$–emittance profile has approximately the same values as its structural connectivity state. For instance, the structural connection from AFDR to the motor neuron RMDVR has weight $\operatorname{{\bf k}}^{AFDR}_{RMDVR}=0.037037$, and its neural emittance onto RMDVR at $\beta_{s}$ is $\operatorname{{\bf x}}^{{AFDR}|\beta_{s}}_{RMDVR}\approx 0.037032$ (see Table 1). Namely, the emittance network of AFDR at inverse temperature $\beta_{s}$ perfectly maps over its anatomical connectivity network.
Diversity in structure-function divergence.
More generally, the $\operatorname{\bf sfd}$ measure provides the ratio of non-overlapping links in both functional states. For instance, $\operatorname{\bf sfd}(AS08,\beta)=0.125$ for $\beta=1.7\beta_{c}$ means that, at this inverse temperature, the emittance network of the motor neuron AS08 deviates from anatomy by 12.5%.
Analyzing $\operatorname{\bf sfd}$ variation of all C. elegans neurons with respect to inverse temperature, we found that the emittance networks of some neurons, such as the motor neurons DD02, AVAL/R, AS04, remain close to their structural connectivity over long intervals, whereas those of other neurons, such as AS08, the polymodal sensory neurons PVDL/R and FLPL/R, and the interneurons LUAL/R, diverge rapidly from structure after certain low temperature values. For example, at $\beta=7.30$, AVAL and AVAR emittance networks deviate from structure by only 1.3%, while PVDL/R deviate by $\sim 9\%$ and AS08 deviates by 12.5% as mentioned earlier.
The c. elegans functional temperature interval.
By computation we found that the mean receptance of C. elegans is 0.5~0.5 at $\beta_{\rm o}=1.07\beta_{c}$ and >0.5>0.5 on the interval $I_{f}=(\beta_{c},\beta_{\rm o})$. For instance, Fig. 1c and Fig. 1d represent the emittance networks of AFDR at two different functional inverse temperatures ($\beta=4.51$ and $\beta=4.34$, respectively).
Purely topological functional connections.
Analyzing the emittance network of AFDR at the structural inverse temperature $\beta_{s}$ shown in Fig. 1b , we found that the emittance onto RMDVR (purple arrow) is not statistically significant ($p\sim.2$, see Table 1), therefore not a PTFC. On the other hand, noting that the anatomical connection from AFDR to RMDVR is defined by a gap-junction, hence a reciprocal connection, we investigate the structural connectivity of RMDVR along with its emittance network (Table 1) at inverse temperature $\beta_{s}$ in order to check whether there is any directional preference on the functioning of this gap-junction. Indeed, we found that the neural emittance of RMDVR onto AFDR is a PTFC ($p=.0056$, see Table 1).
Table 1: Neural emittances of AFDR and RMDVR at $\beta_{s}$.
Note that (in)dependence of a neural emittance connection on topology might vary with (inverse) temperature change. For instance, a PTFC from AFDR to RMDVR eventually emerges significantly ($p=.015$) at a functional temperature $\beta_{f}=1.05\beta_{c}$ (Fig. 1c), albeit with lower weight. This apparent discrepancy is a consequence of the fact that the presence of a large volume of redundant anatomical short paths between two neurons increases the likelihood of these neurons to remain functionally connected at low temperature if the system is randomly rewired, and at higher temperature, redundant long-range pathways increases the likelihood of the emittance networks to be less altered by structural rewiring. For instance, there are $1464$ anatomical paths of length $\leq 3$ all connecting AFDR onto RMDVR, while longer paths between these neurons rely heavily on the existing autapses on RMDVR and other interconnecting neurons such as AFDL, RIAR, IL1DR, etc., forming an infinite number of feedback loops that are very likely destroyed after random reconfiguration of the network.
Figure 2: Emittance network of RID. The pink arrows are connections that are not determined by the network topology, that is, the particular anatomical wiring of the neuron post-synaptic connections, therefore not PTFCs. Green arrows are emittance connections that are topology-dependent, therefore PTFCs. These connections coincide with the functional connections from RID to the neuron classes URX and ADL that have been identified by Randi et al. to be driven by extrasynaptic signaling.
III.3 Putative extrasynaptic signaling pathways
A biological implication of emittance connections that are not PTFCs is that the involved neurons do not necessarily rely on the fixed global anatomy to functionally interact with its target neurons, and additional ’structural’ connections might be needed for their functional connectivity to be predicted from anatomy. This is consistent with results from multiple studies that have demonstrated the presence of extrasynaptic signaling between neurons [48, 49, 50, 51, 52]. To test this interpretation, we investigated the PTFCs of the motor neuron and interneuron RID which has recently been identified by Randi et al. [52] to generate functional connections that were not predicted from anatomy.
On the other hand, considering an extrasynaptic chemical signaling as an actual directed physical connection between neurons, we separately added to the graph one edge from RID to URXL and one edge from RID to ADLR. And by recomputing the $\operatorname{NEP_{\beta}}$ of RID within the updated directed multigraphs, we found that (see Supplementary Data 5 and 6 [39]) the new neural emittances from RID to URXL and ADLR are PTFCs ($p=0.0082$ and $p=0.0086$, respectively).
Figure 3: Purely topological functional network. Schematic of the purely topological functional connectome $\mathcal{F}^{\beta}$ of the C. elegans at a functional inverse temperature value $\beta_{f}=4.41$. It represents only $11.68\%$ of the whole directed weighted network defined by the matrix $\operatorname{{\bf x}}^{\bullet|\beta}]$ whose column vectors are the neural $\beta$-emittance profiles of all neurons. Neurons are positioned according to their spatial coordinates [53].
Figure 4: Purely topological functional connectivity matrix. Representation of part of the adjacency matrix of the purely topological functional connectome $\mathcal{F}^{\beta}$ at $\beta=4.51$. Neurons on the x-axis are the ones with the largest in-degrees in $\mathcal{F}^{\beta}$, and on the y-axis are those with the highest out-degrees.
III.4 Classifying neurons from IC
To investigate the functional properties of the $\operatorname{\bf IC}$, we analyzed its variation with respect to temperature. We observed four main neuron groups showing different integration behaviors (see Table 2 for the complete classification) schematically illustrated by Fig. 5 and characterized as follows.